2. The reaction of bromine gas with chlorine gas, shown here, has a Kc value 0f 7.20 at 200°C. If a close vessel was charged with the two reactants, each at initial concentration of 0.200 M, but with no initial concentration of BrCl, what would be the equilibrium concentration of BrCl?

Br2(g) + Cl2(g) ⇄ 2BrCl(g) Kc=7.20
Initial Concentration 0.200 0.200 0
1
Expert's answer
2020-05-26T14:21:45-0400

Solution:

The balanced chemical equation:

Br2(g) + Cl2(g) = 2BrCl(g)

ICE (Initial, Change, Equilibrium) table for this reaction:




[Br2] = [Cl2] = (0.2 - x) M

[BrCl] = (2x) M

The equilibrium constant expression Kc for the balanced equation:




7.20 = [2x / (0.2 - x)]2

2.6833 = 2x / (0.2-x)

2.6833 × (0.2 - x) = 2x

0.53666 - 2.6833x = 2x

0.53666 = 4.6833x

x = 0.1146


[Br2] = [Cl2] = 0.2 - x = 0.2 - 0.1146 = 0.0854 M

[BrCl] = 2x = (2 × 0.1146) = 0.2292 M

[BrCl] = 0.2292 M


Answer: The equilibrium concentration of BrCl is 0.2292 M.